OpenStudy (anonymous):

Simplify 12k+8k+15j-4k+6j-3j

OpenStudy (anonymous):

\[ 12k +8k +15j -4k +6j -3j =\]

OpenStudy (anonymous):

18j+16k or sqrt(16^2+18^2) at angle = tan-1(16/18)

OpenStudy (anonymous):

16k+18j